# Transfer Learning on Edge Connecting Probability Estimation under Graphon Model

### v1

@@ D.1 Additional results for graph augmentation task
Table D3: Statistics of the datasets used in our experiments.
@@
Table D4: Gromov-Wasserstein distance between each source and target label pair across datasets. Bold values denote the best-matching source label for each target label.
@@
Table D5: Entropic- Gromov-Wasserstein distance with $\epsilon=0.01$ between each source and target label pair across datasets. Bold values denote the best-matching source label for each target label.
@@
D.2 Application to Link Prediction
@@
D.2.1 Transfer from a Single Source Network
Datasets. We evaluate our method on five real-world social networks spanning diverse domains: dolphins, karate, football, firm, and wiki-vote. These datasets cover different interaction types, including animal social behavior (dolphins), human relationships and organizational affiliations (karate, firm), and communication or membership networks (football, wiki-vote). Table D6 summarizes their key statistics. The number of nodes ranges from 33 (firm) to 2,375 (wiki-vote), and the number of edges from 78 (karate) to 16,717 (wiki-vote). The average degree varies accordingly, from 4.59 (karate) to 14.07 (wiki-vote). Network density also differs widely, with firm being relatively dense (0.1723) while wiki-vote is much sparser (0.0074). We select wiki-vote as the source network because it is the largest. Link prediction is then performed on each of the remaining datasets treated as targets, with results averaged over 50 random seeds.
Table D6: Basic statistics of the real-world datasets used in graphon estimation.
@@ Results.
We evaluate transfer performance using Wiki-Vote as the source network ($n_{s}=889$). Each of the remaining datasets (dolphins, firm, football, karate) is treated as the target graph. Results are averaged over 50 random seeds. Table D7 reports the link prediction AUC (mean $\pm$ standard deviation, multiplied by 100). Our method consistently matches or outperforms baselines across datasets. Overall, these results highlight the effectiveness and robustness of transfer-based graphon estimation, particularly in smaller or noisier networks where leveraging external structure is beneficial.
Table D7: Link prediction AUC scores (mean $\pm$ std, $\times 100$). Best result per dataset is bolded.
@@
D.2.2 Intra- vs. Inter-Dataset Transfer
@@
Results. Table D8 reports AUC scores. Transfer from the most structurally similar IMDB-B graph achieves the best performance (AUC 0.96), surpassing both the largest IMDB-B source and all inter-dataset sources. This highlights that transfer is most effective when the source and target share high structural similarity (small GW distance).
Table D8: Link prediction AUC results on target graphs with various transfer sources. For brevity, standard deviations are omitted.
@@
Appendix E Additional Details
E.1 Graphon functions
We implement 10 distinct graphon structures (see Table E9) ranging from simple bilinear forms to highly structured oscillatory and piecewise functions.
Table E9: Graphon functions implemented.
@@ 6 $\frac{xy}{2}$
E.2 Neighborhood Smoothing Details
@@
E.3 Algorithm
E.3.1 Algorithm: GTrans
@@
E.3.2 Algorithm: Network Cross-Validation by Edge Sampling
@@
E.4 Graph Augmentation Model Training Setup.

### v2

@@ F.1 Additional results for graph augmentation task
Table F3: Statistics of the datasets used in our experiments.
@@
Table F4: Gromov-Wasserstein distance between each source and target label pair across datasets. Bold values denote the best-matching source label for each target label.
@@
Table F5: Entropic- Gromov-Wasserstein distance with $\epsilon=0.01$ between each source and target label pair across datasets. Bold values denote the best-matching source label for each target label.
@@
We also include two transfer baselines. (1) Pooled Estimation – pooling source & target graphs and estimating a graphon using the G-Mixup procedure with neighborhood smoothing (NS), and (2) Pooled-then-Transfer – using the pooled estimator as the initial source estimate $\hat{P}^{\text{ini}}_{s}$ in our transfer framework, followed by the GTRANS procedure (denoted PGTRANS-GW & PGTRANS-EGW). From Table F6, both pooled baselines underperform our method, suggesting that naive joint estimation is suboptimal when source and target are misaligned, while our alignment and debiasing steps enable more accurate transfer.
Table F6: Graph classification accuracy (%) comparison between GTRANS and pooled transfer estimation methods (mean $\pm$ std).
F.2 Application to Link Prediction
@@
F.2.1 Transfer from a Single Source Network
Datasets. We evaluate our method on five real-world social networks spanning diverse domains: dolphins, karate, football, firm, and wiki-vote. These datasets cover different interaction types, including animal social behavior (dolphins), human relationships and organizational affiliations (karate, firm), and communication or membership networks (football, wiki-vote). Table F7 summarizes their key statistics. The number of nodes ranges from 33 (firm) to 2,375 (wiki-vote), and the number of edges from 78 (karate) to 16,717 (wiki-vote). The average degree varies accordingly, from 4.59 (karate) to 14.07 (wiki-vote). Network density also differs widely, with firm being relatively dense (0.1723) while wiki-vote is much sparser (0.0074). We select wiki-vote as the source network because it is the largest. Link prediction is then performed on each of the remaining datasets treated as targets, with results averaged over 50 random seeds.
Table F7: Basic statistics of the real-world datasets used in graphon estimation.
@@ Results.
We evaluate transfer performance using Wiki-Vote as the source network ($n_{s}=889$). Each of the remaining datasets (dolphins, firm, football, karate) is treated as the target graph. Results are averaged over 50 random seeds. Table F8 reports the link prediction AUC (mean $\pm$ standard deviation, multiplied by 100). Our method consistently matches or outperforms baselines across datasets. Overall, these results highlight the effectiveness and robustness of transfer-based graphon estimation, particularly in smaller or noisier networks where leveraging external structure is beneficial.
Table F8: Link prediction AUC scores (mean $\pm$ std, $\times 100$). Best result per dataset is bolded.
@@
F.2.2 Intra- vs. Inter-Dataset Transfer
@@
Results. Table F9 reports AUC scores. Transfer from the most structurally similar IMDB-B graph achieves the best performance (AUC 0.96), surpassing both the largest IMDB-B source and all inter-dataset sources. This highlights that transfer is most effective when the source and target share high structural similarity (small GW distance).
Table F9: Link prediction AUC results on target graphs with various transfer sources. For brevity, standard deviations are omitted.
@@
Appendix G Additional Details
G.1 Graphon functions
We implement 10 distinct graphon structures (see Table G10) ranging from simple bilinear forms to highly structured oscillatory and piecewise functions.
Table G10: Graphon functions implemented.
@@ 6 $\frac{xy}{2}$
G.2 Neighborhood Smoothing Details
@@
G.3 Algorithm
G.3.1 Algorithm: GTrans
@@
G.3.2 Algorithm: Network Cross-Validation by Edge Sampling
@@
G.4 Graph Augmentation Model Training Setup.